# Deep-learning-based continuous attacks on QKD protocols

### v3

@@ Appendix G Information gain computation
[…]
Let $S$ be a discrete variable representing the different initial states ($s=0,1,2,3$) and $X\in\mathbb{R}^{40}$ a $40$-dimensional continuous variable representing the values of the homodyne currents at each time step. The joint distribution can be written $P_{X,S}(x,s)=P(s)P(x|s)$, such that Eq. (55) becomes
[…]
@@
To compute Eq. (56), we assume that for each state $s$, the measurement vectors follow a multivariate Gaussian distribution
$P(x\mid s)=\mathcal{N}\left(x\mid\mu_{s},\Sigma_{s}\right).$ (58)
We test this assumption with the Henze-Zirkler multivariate normality test (over $2\times 10^{4}$ samples), which tests the null hypothesis that the data follows a multivariate normal distribution [98]. The results are displayed in Table 3.
State p-value
|$\ket{0}$ $0.8757$
|$\ket{1}$ $0.7008$
$\ket{+}$ $0.9424$
$\ket{-}$ $0.7862$
Table 3: P-values for the Henze-Zirkler multivariate normality test. A p-value above $0.05$ suggests the data could plausibly follow a multivariate normal distribution.
Conditional distributions are shown for each state in Fig. 11.
[…]
The logarithms in these formulae are usually taken to base 2 if information is measured in units of bits, or to base e if information is measured in nats. The resulting information gain is
$I(S;x)$ $=0.2894\ \text{nats}=0.4175\ \text{bits}.$

### v4

@@ Appendix E Information gain computation
$I(X;Y)=\iint P_{X,Y}(x,y)\log\left(\frac{P_{X,Y}(x,y)}{P_{X}(x)P_{Y}(y)}\right)dx\,dy,$ (51)
which quantifies how much knowing $X$ reduces uncertainty about $Y$.
Let $S$ be a discrete variable representing the different initial states ($s=0,1,2,3$) and $X\in\mathbb{R}^{70}$ a $70$-dimensional continuous variable representing the values of the homodyne currents at each time step. The joint distribution can be written $P_{X,S}(x,s)=P(s)P(x|s)$, such that Eq. (51) becomes
$I(X;S)=\sum_{s}P(s)\int P(x|s)\log\left(\frac{P(x|s)}{P(x)}\right)dx.$ (52)
@@
To estimate Eq. (52), we used a non-parametric approach based on entropy estimates from k-nearest neighbor distances, which was first presented by Kraskov et al. in [96]. For such numerical estimation, the Python package NPEET-plus (Non Parametric Entropy Estimation Toolbox) provides confidence interval estimation for mixed mutual information using bootstrapping. The currents were randomly sub-sampled to a $3\times 10^{4}$ subset, and (per-feature-)standardized to zero mean and unit variance. Their dimensionality was then reduced, using Principal Component Analysis, to a single component, which served for the mutual information estimation.
Mutual information between model predictions and ground truth
Since the predictions and the ground truth are discrete variables, respectively denoted $\hat{y}$ and $y$, Eq. (51) becomes
$I(\hat{Y};Y)=\underset{\hat{y},y}{\sum}P(\hat{y},y)\log\left(\frac{P(\hat{y},y)}{P(\hat{y})P(y)}\right),$ (54)
which we compute from the confusion matrix of the model, displayed in Fig. 7.
Figure 7: Confusion matrix for our 4-class classification problem. Rows are ground truth while columns are predictions
The confusion matrix $(i,j)$ entries are the empirical joint probabilities $P(\hat{Y}_{i},Y_{j})$, which we approximate the true probabilities with, and the marginal distributions P(Y^)P($\hat${Y}) and $P(Y)$ are the sum of the rows and the columns respectively. Plugging this into Eq. (54), we obtain
$I(\hat{Y};Y)=0.1527\ \text{bits}.$ (55)